OpenStudy (anonymous):

does anyone know a function with two irrational solutions, and another with two complex solutions?

OpenStudy (anonymous):

you can create your self let the roots be \[\sqrt{2},\sqrt{3},2\iota,-2\iota \] then equation is \[\left( x-\sqrt{2} \right)\left( x-\sqrt{3} \right)\left( x-2\iota \right)\left( x+2\sqrt{\iota } \right)=0\] simplify it.

OpenStudy (radar):

f(x) = x^2 + 1 would give you two complex roots. (x + i) (x-i) =0 x= -i x= i those two could be considered complex with 0 real values.

OpenStudy (anonymous):

\[f(x)=x^2-5x+1\] take any quadratic whose determinant >0 and is not a perfect square.